OpenStudy (anonymous):

Show all necessary work for the following problems. All answers must be in SIMPLEST radical form and fractions in lowest terms. 1.√48 2.√135 3.√48x^4 y^5 4.12/√3 5.8√32/2√8 6.√72x^9 y^4 z^6/√6x^4 y^5 7.7√2/5+√6 8.8/2-√10 9.2√5 (3√14+3√7) 10.(4√6-4)^2 11.3√49+5√52-√18 12.4√-x=24 13.√x+7-2=x 14. 4√10x/5+7=15 15. √x-6=x-8 16. Find the perimeter of a rectangle whose length is 3√3+7 and whose width is √6-8.

OpenStudy (anonymous):

1. sqrt(16*3) =sqrt(3) 4. You just find perfect squares on the inside and then you can pull them out. 2.sqrt(15*9)= sqrt(15) 9=sqrt(3)sqrt(5)9. ext ext...

OpenStudy (anonymous):

holy moses!!!

OpenStudy (anonymous):

Can anyone please answer all these problems? If you do, I'll appreciate it.

OpenStudy (anonymous):

http://www.freemathhelp.com/ go to the algebra tab and sign up for a free 1 day trial with your email address and it will show all the steps it takes to getting the answer. (:

OpenStudy (anonymous):

Thanks but I already tried that before so that won't help.

OpenStudy (anonymous):

use a diff email. then